OpenStudy (anonymous):

solve the system, using elimination 2x+6y=-12, 5x-5y=10 a. (2,1) b. (0,-2) c. (-2,0) d. (1,2)

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y=-x/3-2, y=x-2 Plug in x for y and you shall see it is b.